2. 166 points E2-6 Finding Unknown Values in the Cost of Goods Manufactured Report ILO 2-3, 2-6) Muligan Manufacturing Company uses a job order cost system with ovarhead applied to products at a rate of 150 percent of direct labor cost. Required: Treating each case independently, selected from the manufacturing data given below, find the missing amounts. You should do them in the order listed. (Hint: For the e manufacting costs in Case 3, first solve for conversion costs and then determine how much of that is dwect labor and how much is manufactuning overhead) (Do not round your intermediate calculations. Round your final answers to the nearest whole dollar. Enter all amounts as positive values.) Case1 Case 2 Case 3 Direct material used Direct labor Manufacturing overhead applied Total current manufacturing costs Beginning work in process inventory Ending work in process inventory Cost of goods manutactured Beginning firished goods inventory Ending finished goods inventory Cost of goods sold 12,000 9,800 12,000 12,500 29,000 28,200 8,100 7,200 8,900 43,000 10.400 10,001 8,900 6,600 42,000 37,000
Answer #1

Answer:

Determine missing amount :
Caln Case 1($) Case 2($) Case 3($)
direct material used A                              12,000                      8,167                      9,600
Direct labour           B                              12,000                      8,333                      7,440
Manufacturing overhead applied(150% of Cirect labour)      C                              18,000                    12,500                   11,160
total current manufacturing costs D=A+B+C                              42,000                    29,000                    28,200
beginning finished work in process inventory E                                 8,100                   22,900                      7,200
ending work in process inventory(less) F                                 5,900                      8,900                   25,399
cost of goods manufactured G=D+E-F                              44,200                    43,000                    10,001
beginning finished goods inventory H                                 4,900                    10,400                   38,599
ending finished goods inventory I                                 8,900                   16,400                      6,600
Cost of goods sold J=G+H-I                              40,200                    37,000                    42,000
Case 3 -Direct labour =28200-9600/2.5= $7,440